# Closing the Gap Between SGP4 and High-Precision Propagation via Differentiable Programming

### 27 Feb 2024 v3

@@ Abstract
The simplified general perturbations 4 (SGP4) orbital propagation is one of the most widely used approaches used to compute a rapid and relatively reliable prediction of the positions and velocity of objects orbiting Earth. Over time, SGP models have undergone refinement to enhance their efficiency and accuracy. Yet, they still fall short of the precision offered by numerical propagators, which can predict the positions and velocities of space objects in low-Earth orbit with significantly smaller errors. In this study, we introduce a novel differentiable version of SGP4, named dSGP4. By reconfiguring SGP4 into a differentiable program using PyTorch, we unlock its differentiability for various space-related applications, encompassing spacecraft orbit determination, state conversion, covariance similarity transformation, state transition matrix computation, covariance propagation, and more. A side advantage of this PyTorch-based model is its capacity for "embarrassingly parallel" orbital propagation across batches of TLEs. It can harness the processing power of CPUs, GPUs, and advanced hardware alike, to be able to predict batches of satellites at future times in a distributed fashion. Furthermore, the inherent differentiability of dSGP4 renders it compatible with modern machine learning methodologies. Consequently, we propose a novel orbital propagation paradigm, ML-dSGP4. In this paradigm, the orbital propagator is imbued with neural networks. Through stochastic gradient descent during the learning process, the inputs, outputs, and parameters of this combined model can be iteratively refined to achieve precision surpassing that of SGP4. Fundamentally, the neural networks function as identity operators when the propagator adheres to its default behavior as defined by SGP4. However, owing to the differentiability ingrained within dSGP4, the model can be fine-tuned with ephemeris data to learn corrections to both inputs and outputs of SGP4. This augmentation enhances precision while maintaining the same computational speed of dSGP4 at inference time. This empowers satellite operators and researchers, equipping them with the ability to train the model using their specific ephemeris or high-precision numerical propagation data.
@@ 1 Introduction
Motivated by these works, and by the lack of a general framework that allows improving SGP4 predictions integrating highly accurate data and/or predictions, we developed a differentiable SGP4 (dSGP4) version on Python, which we make available through an open-source repository on the European Space Agency Github organization 333https://github.com/esa/dSGP4, date of access: March 2023.. Our contributions consist of first re-writing the SGP4 propagator to make it compatible with modern programming techniques, which enable automatic differentiation and "embarrassingly parallel" computations via CPU and/or GPU. We then show that having an SGP4 propagator that is differentiable with respect to its inputs opens the possibility of using the Jacobian for several applications (e.g. state covariance propagation, state transition matrix computation, gradient-based optimization). This has also consequences in many fields related to SSA, such as orbit determination, among others. To underscore the significance of employing dSGP4 in practical contexts, we reference a work by Vallado on SGP4 [21], where he suggests that while elegant operational systems typically opt for numerical integration, a straightforward analytical propagation theory like SGP4 might justify employing finite differencing. According to him, the effort and resources invested in developing partial derivatives for various data types and state forms to achieve computational efficiency are often outweighed by the simplicity of incorporating an additional iteration or two with a finite differencing approach. Through the use of dSGP4, we demonstrate that the computation of partial derivatives can be seamlessly executed without incurring any extra time or financial costs, thanks to automatic differentiation. Additionally, this approach mitigates local truncation errors associated with finite differencing.
@@
Finally, we also discuss the applications of such a differentiable SGP4 model in a new hybrid propagator paradigm, where the machine learning model is made of an initial neural network that ingests the data, which is then passed through dSGP4, and finally through another neural network. In this way, owing to the differentiability of dSGP4, it is possible to take gradients of the entire model, therefore applying standard stochastic gradient descent techniques and training dSGP4 together with neural network models. This framework enables backpropagating through the whole model, and learning corrections of SGP4 inputs and outputs to match the predicted state: in practice, this means that the neural networks, if successful, learn the differential correction terms needed to correct the low-accuracy predictions of SGP4, to better match either observations or high-fidelity propagations, and retain the advantageous SGP4 speed. In this way, for the first time, both inputs and outputs (and potentially hyperparameters of SGP4) can be learned for the given prediction task, to maximize the accuracy of the predicted orbital states, while maintaining a fast and parallelizable propagator.
@@
The paper is organized as follows: in Section 2, we discuss relevant background on the SGP4 model and differentiable programming. Then, in Section 3, we discuss the features of the propagator, including automatic differentiation and batch mode. In Sec. 4, we discuss its machine learning integration, by both outlining the general framework and discussing simulation results that include the use of the machine learning dSGP4 framework to improve the SGP4 predictions on a batch of SpaceX Starlink satellites. While in Sec. 5, we introduce the use of dSGP4 in the context of satellite orbit determination. Other applications including gradient-based optimization, osculating to TLE elements conversion, covariance propagation, similarity transformation, state transition matrix computation, and batch propagation are included as tutorials of the open-source library and released with dSGP4. Finally, we also conduct an experiment on a group of SpaceX Starlink satellites where dSGP4 and ephemeris data are used to train the ML-enhanced hybrid propagator, to improve the baseline SGP4 predictions. Finally, in Section 6, we highlight the conclusions of our work.
@@ 2.2 Differentiable Programming
Differentiable programming is a programming paradigm in which models can be built through a computer program and then their parameters can be differentiated through automatic differentiation (AD). This allows users to perform gradient-based optimization of the model outputs with respect to its parameters. This means that users can leverage gradient-based techniques to optimize large and complex models with many parameters, adjusting those parameters in order to minimize the deviation of the model output with respect to ground truth data. With the advent of artificial intelligence and neural network models, this paradigm has become increasingly popular, and many scientific fields have successfully used differentiable programming in several areas: from computer science to robotics, ray tracing, image processing, probabilistic programming, electronics, etc. [22, 23, 24, 25].
By investigating the SGP4 formulation, it is apparent that it can be described as a set of continuous and differentiable functions that given several input parameters, returns a 6-dimensional vector, and that this can be used in some applications such as orbit determination [26, 4]. However, there is currently no open-source SGP4 version formulated as a differentiable program, on which automatic differentiation can be applied not only for direct gradient-based optimization techniques such as orbit determination and model calibration but also to combine deep learning with SGP4. In fact, one of the current issues of the application of SGP4 in SSA is that while it allows performing fast orbital prediction and conjunction assessment for the full catalog, its results are not particularly accurate, which forces many operators to refine their analysis with more precise (but slower) numerical integration schemes that represent the underlying dynamical model more accurately [21, 16, 27]. Some researchers have attempted to trade off the SGP4 speed with numerical integration accuracy [28] or to take advantage of neural networks to correct the output of SGP4 [7, 19]. However, to our knowledge, there have been no attempts to combine SGP4 and neural networks within the same model to maintain the same speed as SGP4 while correcting the model’s inputs, outputs, and hyperparameters to better match accurate data or high-fidelity orbital propagators. In this work, we explore this aspect and show that formulating SGP4 as a differentiable program paves the way to easy integration with neural networks.
3 dSGP4: an Open-Source Differentiable SGP4 Program
3.1 dSGP4 Overview: Automatic Differentiation and Batch Mode
@@
Automatic differentiation is a technique used for computing and evaluating derivatives of functions. It can be applied to any function that can be represented as a sequence of elementary operations, whose derivatives are known [29]. The basic idea of AD is to build a graph of elementary operations with known derivatives, which is then used (either in forward or reverse mode) to precisely compute the derivative at a given point. Before automatic differentiation, derivatives could either be computed manually, with a time-consuming and error-prone procedure, through symbolic differentiation, or through numerical differentiation. This last alternative is not ideal due to the presence of truncation and round-off errors, which might substantially jeopardize the accuracy of the found derivatives, if not chosen carefully. On the other hand, symbolic differentiation would allow to access the derivative of the function for the entire domain, but it can result in the size of the expressions growing exponentially (something known as "expression swell"), as the number of variables and operations is increased, due to the product rule of derivative [30]. Automatic differentiation solves these issues, by returning the exact value (up to machine precision) of the derivatives of a computer program evaluated at a certain point, without incurring exponentially growing expressions [31]. The drawback, besides the fact that it is not applicable for non-differentiable functions, is that this is only the derivative evaluated at a given point, and one does not have access to the symbolic form of the derivative across the entire domain. The differentiability constraint implies that the computer program needs to comply with certain requirements and does not admit discontinuities, singularities, or non-differentiable features. In our case, we have implemented dSGP4 using the Python library Pytorch [32], which is one of the most popular open-source ML libraries that supports AD in use in deep learning research and applications. To ensure the fidelity of dSGP4, our approach encompassed several key verification steps. First of all, we validated the dSGP4 model’s outputs against the "Revisiting Spacetrack Report #3" SGP4 model [10]. This involved the design and execution of continuous integration tests (that can be found in the open-source library) to confirm that all mirrored functionalities faithfully reproduced the behavior of the original implementation444https://github.com/esa/dSGP4/tree/master/tests, date of access: February 2024.. As a part of these tests, we also verified that we could reproduce the output of the Space Track official SGP4 release555https://www.space-track.org/documentation#/sgp4, date of access: February 2024.: this was done by propagating several different TLEs at different future times (between a few hours to a few days), and checking that the results of the propagation match up to below 1 mm in position and 1 mm/s in velocity the official release results. Our verifications also extended to the examination of automatic differentiation outputs for the input variables. To ensure correctness, we compared (up to local truncation errors caused by the finite difference technique) that the results obtained from dSGP4 automatic differentiation functionality match the ones found via finite differencing of the original SGP4.
@@
As previously discussed, dSGP4 also offers an additional capability, as we show in Figure 1. While retaining the traditional functionality of predicting object states from a TLE for specific times or multiple times, dSGP4 offers a distinctive feature called batch mode. In this mode, batches of TLEs can be provided to the model, which in turn returns the corresponding batch of Cartesian states. This expanded functionality has also undergone comprehensive testing within our library.
@@
Differentiable physics simulators have the key advantage that can be run both in forward and backward mode. When the forward mode is active, the simulator has the commonly known behavior of returning a certain output from the provided inputs: however, when the backward mode is activated, the model can also return the gradients of the outputs with respect to any input and parameter of the model. This allows the integration of this simulator with gradient-based methods, where the gradient information can be leveraged to learn and control the physical system. This approach enables a variety of inverse problems that involve orbital propagation, which we will also discuss in the coming sections (e.g. satellite orbit determination). The interested reader can also find in the dSGP4 website 666https://esa.github.io/dSGP4, date of access: October 2023. a plethora of tutorials on how to apply the dSGP4 framework to common gradient-based tasks for SSA applications (e.g. state transition matrix and sensitivity matrix computation, covariance propagation, and gradient-based optimization). The simulator is treated as a differentiable function, and a certain objective function is minimized, which depends on the simulator. Mathematically this is formulated as a minimization problem of a certain objective function: , where is an -dimensional vector of model parameters. The goal is to update the parameters in such a way that reaches the minimum value: this is done in the direction opposite of the gradient of the objective function w.r.t. the parameters. In its simplest (i.e., vanilla gradient descent) form this can be written as , where is known as learning rate [33]. Therefore, it is clear that by formulating the orbital propagator as a differential program, its integration with any gradient-related task, such as vanilla gradient descent is possible. This also has the direct consequence of being able to use these gradients within root-finding algorithms that leverage gradient information.
@@ 3.2 Performance Profiling
As already discussed in the previous section, an important advantage of having a differentiable SGP4 program written in PyTorch is the possibility to massively distribute the propagation of TLEs, which we refer to as batch-mode propagation, leveraging either multiple CPUs or GPU acceleration (through CUDA). In this way, we enable the possibility to have embarrassingly parallel computations of multiple TLEs at multiple times, through hardware parallelization. To quantitatively assess this capability, we display in Tab. 1 the performance of the simple SGP4, against dSGP4 on 1 CPU and GPU. We compare the results using float64 numbers (this is not ideal for the GPU case, but it might be important for orbital propagation applications). In terms of CPU processor, we have used the MacBook M1 Pro processor, while in terms of GPU, we have used the NVIDIA V100 Tensor Core GPU777https://www.nvidia.com/en-us/data-center/v100/, date of access: August 2023.. As we see from the table, already for batches of size 1,000, dSGP4 offers a significant speed-up, when run in batch mode, on both CPU and GPU, compared to the standard SGP4 program. In particular, as we see, the GPU implementation can reach a factor of 85 speed-up compared to the standard SGP4. This level of speed-up factor holds promise for applications such as all-versus-all conjunction screening, where computational speed-up can yield substantial benefits.
@@
dSGP4 (1 CPU) | 0.67 | 4.52 | 44.6 | 460 | 7,920 |
dSGP4 (GPU) | 4.22 | 7.8 | 8 | 29 | 249 |
4 Machine Learning dSGP4
@@ 4.1 Methods
This section introduces ML-dSGP4, a paradigm where a differentiable propagator (dSGP4) is integrated with two neural networks to correct its inputs and outputs. In Figure 1(a) and 1(b), respectively, both standard SGP4 and dSGP4 propagations are illustrated. While in Figure 1(b) and 1(d), respectively, the combination between ML and both SGP4 and dSGP4 are shown.
The ML-dSGP4 model represents a new hybrid differentiable propagator that retains the computational advantages of SGP4 while enabling training and improvement with accurate data. The model can be configured to return either SGP4 predictions or activate input and output corrections using trained neural networks.
To better discuss the ML-dSGP4 setup shown in Figure 2, let us call the input of ML-dSGP4 (i.e., the TLE elements), the propagation time, and its output, and let us assume that the objective is to minimize the error between the ML-dSGP4 propagation and the propagation of a high precision propagator, (i.e., a numerical integrator that accounts for all forces involved in the satellite motion). In this case, the objective is to correct ML-dSGP4 to better match the accurate propagator, hence, the function to minimize will be a function of the state, time, and parameters (both of the networks and also of dSGP4). Without loss of generality, let’s consider as an objective function the squared norm between the target state and the dSGP4 output, then the objective is to find the parameters (i.e., weights and biases) of the two networks, as well as the parameters of SGP4 that minimize the loss:
where indicates the parameters of the input neural network, of the output one, and the SGP4 parameters. Owing to the differentiability of the neural networks and dSGP4, we can use gradient descent to find the local minimum of the loss function by iteratively updating the parameters of the neural networks and dSGP4:
This setup can be easily extended for multiple propagation points, for instance, by taking the mean squared error as a loss function, and by setting up a stochastic gradient descent algorithm to update the parameters. The advantage of ML-dSGP4 lies in its capacity to learn both the initial neural network weights and SGP4 parameters, therefore offering a flexible solution. This differs from ML-SGP4, where only the output of the SGP4 propagator is corrected, as explored in prior research [7]. In the next section, we will delve into some numerical experiments to showcase the use of ML-dSGP4 and to highlight the comparison between ML-dSGP4 and the ML-SGP4.

### 07 Mar 2024 v4

@@ Abstract
The simplified general perturbations 4 (SGP4) orbital propagation is one of the most widely used approaches for computing a rapid and relatively reliable prediction of the positions and velocity of objects orbiting Earth. Over time, SGP models have undergone refinement to enhance their efficiency and accuracy. Yet, they still fall short of the precision offered by high-precision numerical propagators, which can predict the positions and velocities of space objects in low-Earth orbit with significantly smaller errors. In this study, we introduce a novel differentiable version of SGP4, named SGP4. By reconfiguring SGP4 into a differentiable program using PyTorch, we unlock its differentiability for various space-related applications, encompassing spacecraft orbit determination, state conversion, covariance similarity transformation, state transition matrix computation, covariance propagation, and more. A side advantage of this PyTorch-based model is its capacity for "embarrassingly parallel" orbital propagation across batches of TLEs. It can harness the processing power of CPUs, GPUs, and advanced hardware alike, to be able to predict batches of satellites at future times in a distributed fashion. Furthermore, the inherent differentiability of SGP4 renders it compatible with modern machine learning methodologies. Consequently, we propose a novel orbital propagation paradigm, ML-SGP4. In this paradigm, the orbital propagator is imbued with neural networks. Through stochastic gradient descent during the learning process, the inputs, outputs, and parameters of this combined model can be iteratively refined to achieve precision surpassing that of SGP4. Fundamentally, the neural networks function as identity operators when the propagator adheres to its default behavior as defined by SGP4. However, owing to the differentiability ingrained within SGP4, the model can be fine-tuned with ephemeris data to learn corrections to both inputs and outputs of SGP4. This augmentation enhances precision while maintaining the same computational speed of SGP4 at inference time. This empowers satellite operators and researchers, equipping them with the ability to train the model using their specific ephemeris or high-precision numerical propagation data.
@@ 1 Introduction
Motivated by these works, and by the lack of a general framework that allows improving SGP4 predictions integrating highly accurate data and/or predictions, we developed a differentiable SGP4 (SGP4) version on Python, which we make available through an open-source repository on the European Space Agency Github organization333https://github.com/esa/dSGP4, date of access: March 2023.. Our contributions consist of first re-writing the SGP4 propagator to make it compatible with modern programming techniques, which enable automatic differentiation and "embarrassingly parallel" computations via CPU and/or GPU. We then show that having an SGP4 propagator that is differentiable with respect to its inputs opens the possibility of using the Jacobian matrix for several applications (e.g. state covariance propagation, state transition matrix computation, gradient-based optimization). This has also consequences in many fields related to SSA, such as orbit determination, among others. To underscore the significance of employing SGP4 in practical contexts, we reference a work by Vallado on SGP4 [21], where he suggests that while elegant operational systems typically opt for numerical integration, a straightforward analytical propagation theory like SGP4 might justify employing finite differencing. According to him, the effort and resources invested in developing partial derivatives for various data types and state forms to achieve computational efficiency are often outweighed by the simplicity of incorporating an additional iteration or two with a finite differencing approach. Through the use of SGP4, we demonstrate that the computation of partial derivatives can be seamlessly executed without incurring any extra time or financial costs, thanks to automatic differentiation. Additionally, this approach mitigates local truncation errors associated with finite differencing.
@@
Finally, we also discuss the applications of such a differentiable SGP4 model in a new hybrid propagator paradigm, where the machine learning model is made of an initial neural network that ingests the data, which is then passed through SGP4, and finally through another neural network. In this way, owing to the differentiability of SGP4, it is possible to take gradients of the entire model, therefore applying standard stochastic gradient descent techniques and training SGP4 together with neural network models. This framework enables backpropagating through the whole model, and learning corrections of SGP4 inputs and outputs to match the predicted state: in practice, this means that the neural networks, if successful, learn the differential correction terms needed to correct the low-accuracy predictions of SGP4, to better match either observations or high-fidelity propagations, and retain the advantageous SGP4 speed. Thus, for the first time, both inputs and outputs (and potentially hyperparameters of SGP4) can be learned for the given prediction task, to maximize the accuracy of the predicted orbital states, while maintaining a fast and parallelizable propagator.
@@
The paper is organized as follows: in Section 2, we discuss relevant background on the SGP4 model and differentiable programming. Then, in Section 3, we discuss the features of the propagator, including automatic differentiation and batch mode. In Sec. 4, we discuss its machine learning integration, by both outlining the general framework and discussing simulation results that include the use of the machine learning SGP4 framework to improve the SGP4 predictions on a batch of SpaceX Starlink satellites. While in Sec. 5, we introduce the use of SGP4 in the context of satellite orbit determination. Other applications including gradient-based optimization, osculating to TLE elements conversion, covariance propagation, similarity transformation, state transition matrix computation, and batch propagation are included as tutorials of the open-source library and released with SGP4. Then, we also conduct an experiment on a group of SpaceX Starlink satellites where SGP4 and ephemeris data are used to train the ML-enhanced hybrid propagator, to improve the baseline SGP4 predictions. Finally, in Section 6, we highlight the conclusions of our work.
@@ 2.2 Differentiable Programming
Differentiable programming is a programming paradigm in which models can be built through a computer program and then their parameters can be differentiated through automatic differentiation (AD). This allows users to perform gradient-based optimization of the model outputs with respect to its parameters. This means that users can leverage gradient-based techniques to optimize large and complex models with many parameters, adjusting those parameters in order to minimize the deviation of the model output with respect to ground truth data. With the advent of artificial intelligence and neural network models, this paradigm has become increasingly popular, and many scientific fields have successfully used differentiable programming in several areas: from computer science to robotics, ray tracing, image processing, probabilistic programming, electronics, etc. [22, 23, 24, 25]
3 SGP4: an Open-Source Differentiable SGP4 Program
3.1 SGP4 Overview: Automatic Differentiation and Batch Mode
@@
Automatic differentiation is a technique used for computing and evaluating derivatives of functions. It can be applied to any function that can be represented as a sequence of elementary operations, whose derivatives are known [26]. The basic idea of AD is to build a graph of elementary operations with known derivatives, which is then used (either in forward or reverse mode) to precisely compute the derivative at a given point. Before automatic differentiation, derivatives could either be computed manually, with a time-consuming and error-prone procedure, through symbolic differentiation, or through numerical differentiation. This last alternative is not ideal due to the presence of truncation and round-off errors, which might substantially jeopardize the accuracy of the found derivatives, if not chosen carefully. On the other hand, symbolic differentiation would allow to access the derivative of the function for the entire domain, but it can result in the size of the expressions growing exponentially (something known as "expression swell"), as the number of variables and operations is increased, due to the product rule of derivative [27]. Automatic differentiation solves these issues, by returning the exact value (up to machine precision) of the derivatives of a computer program evaluated at a certain point, without incurring exponentially growing expressions [28]. The drawback, besides the fact that it is not applicable for non-differentiable functions, is that this is only the derivative evaluated at a given point, and one does not have access to the symbolic form of the derivative across the entire domain. The differentiability constraint implies that the computer program needs to comply with certain requirements and does not admit discontinuities, singularities, or non-differentiable features. In this case, we have implemented SGP4 using the Python library Pytorch [29], which is one of the most popular open-source ML libraries that supports AD in use in deep learning research and applications. To ensure the fidelity of SGP4, our approach encompassed several key verification steps. First of all, we validated the SGP4 model’s outputs against the "Revisiting Spacetrack Report #3" SGP4 model [10]. This involved the design and execution of continuous integration tests (that can be found in the open-source library) to confirm that all mirrored functionalities faithfully reproduced the behavior of the original implementation444https://github.com/esa/dSGP4/tree/master/tests, date of access: February 2024.. As a part of these tests, we also verified that we could reproduce the output of the Space Track official SGP4 release555https://www.space-track.org/documentation#/sgp4, date of access: February 2024.: this was done by propagating several different TLEs at different future times (between a few hours to a few days), and checking that the results of the propagation match up to below 1 mm in position and 1 mm/s in velocity the official release results. Our verifications also extended to the examination of automatic differentiation outputs for the input variables. To ensure correctness, we compared (up to local truncation errors caused by the finite difference technique) that the results obtained from SGP4 automatic differentiation functionality match the ones found via finite differencing of the original SGP4.
@@
As previously discussed, SGP4 also offers an additional capability, as we show in Figure 1. While retaining the traditional functionality of predicting object states from a TLE for specific times or multiple times, SGP4 offers a distinctive feature called batch mode. In this mode, batches of TLEs can be provided to the model, which in turn returns the corresponding batch of Cartesian states. This expanded functionality has also undergone comprehensive testing within our library.
@@
Differentiable physics simulators have the key advantage that can be run both in forward and backward mode. When the forward mode is active, the simulator has the commonly known behavior of returning a certain output from the provided inputs: however, when the backward mode is activated, the model can also return the gradients of the outputs with respect to any input and parameter of the model. This allows the integration of this simulator with gradient-based methods, where the gradient information can be leveraged to learn and control the physical system. This approach enables a variety of inverse problems that involve orbital propagation, which we will also discuss in the coming sections (e.g. satellite orbit determination). The interested reader can also find in the SGP4 website 666https://esa.github.io/$∂$SGP4, date of access: October 2023. a plethora of tutorials on how to apply the SGP4 framework to common gradient-based tasks for SSA applications (e.g. state transition matrix and sensitivity matrix computation, covariance propagation, and gradient-based optimization). The simulator is treated as a differentiable function, and a certain objective function is minimized, which depends on the simulator. Mathematically this is formulated as a minimization problem of a certain objective function: , where is an -dimensional vector of model parameters. The goal is to update the parameters in such a way that reaches the minimum value: this is done in the direction opposite of the gradient of the objective function w.r.t. the parameters. In its simplest (i.e., vanilla gradient descent) form this can be written as , where is known as learning rate [30]. Therefore, it is clear that by formulating the orbital propagator as a differential program, its integration with any gradient-related task, such as vanilla gradient descent is possible. This also has the direct consequence of being able to use these gradients within root-finding algorithms that leverage gradient information.
@@ 3.2 Performance Profiling
As already discussed in the previous section, an important advantage of having a differentiable SGP4 program written in PyTorch is the possibility to massively distribute the propagation of TLEs, which we refer to as batch-mode propagation, leveraging either multiple CPUs or GPU acceleration (through CUDA). In this way, we enable the possibility to have embarrassingly parallel computations of multiple TLEs at multiple times, through hardware parallelization. To quantitatively assess this capability, we display in Tab. 1 the performance of the simple SGP4, against SGP4 on 1 CPU and GPU. We compare the results using float64 numbers (this is not ideal for the GPU case, but it might be important for orbital propagation applications). In terms of CPU processor, we have used the MacBook M1 Pro processor, while in terms of GPU, we have used the NVIDIA V100 Tensor Core GPU777https://www.nvidia.com/en-us/data-center/v100/, date of access: August 2023.. As we see from the table, already for batches of size 1,000, SGP4 offers a significant speed-up, when run in batch mode, on both CPU and GPU, compared to the standard SGP4 program. In particular, as we see, the GPU implementation can reach a factor of 85 speed-up compared to the standard SGP4. This level of speed-up factor holds promise for applications such as all-versus-all conjunction screening, where computational speed-up can yield substantial benefits.
@@
SGP4 (1 CPU) | 0.67 | 4.52 | 44.6 | 460 | 7,920 |
SGP4 (GPU) | 4.22 | 7.8 | 8 | 29 | 249 |
4 Machine Learning SGP4
@@ 4.1 Methods
This section introduces ML-SGP4, a paradigm where a differentiable propagator (SGP4) is integrated with two neural networks to correct its inputs and outputs. In Figure 1(a) and 1(b), respectively, both standard SGP4 and SGP4 propagations are illustrated. While in Figure 1(b) and 1(d), respectively, the combination between ML and both SGP4 and SGP4 are shown.
The ML-SGP4 model represents a new hybrid differentiable propagator that retains the computational advantages of SGP4 while enabling training and improvement with accurate simulated or observed data. The model can be configured to return either SGP4 predictions or activate input and output corrections using trained neural networks.
To better discuss the ML-SGP4 setup shown in Figure 2, let us call the input of ML-SGP4 (i.e., the TLE elements), the propagation time, and its output, and let us assume that the objective is to minimize the error between the ML-SGP4 propagation and the propagation of a high precision propagator, (i.e., a numerical integrator that accounts for all forces involved in the satellite motion). In this case, the objective is to correct ML-SGP4 to better match the accurate propagator, hence, the function to minimize will be a function of the state, time, and parameters (both of the networks and also of SGP4). Without loss of generality, let’s consider as an objective function the squared norm between the target state and the SGP4 output, then the objective is to find the parameters (i.e., weights and biases) of the two networks, as well as the parameters of SGP4 that minimize the loss:
where indicates the parameters of the input neural network, of the output one, and the SGP4 parameters. Owing to the differentiability of the neural networks and SGP4, we can use gradient descent to find the local minimum of the loss function by iteratively updating the parameters of the neural networks and SGP4:
This setup can be easily extended for multiple propagation points, for instance, by taking the mean squared error as a loss function, and by setting up a stochastic gradient descent algorithm to update the parameters. A schematic illustration of this architecture is shown in Figure 3. As it can be seen, the model is first run in forward mode: in this way, from a set of TLEs, a set of propagated states can be found. Then, in backward mode, the loss between these states and simulated (e.g. HPOP simulations) or observed states (e.g. ephemeris data) is computed via automatic differentiation, and the partials of the loss w.r.t. both neural network parameters and SGP4 can be used to update the models, until convergence.
The advantage of ML-SGP4 lies in its capacity to learn both the neural network weights and SGP4 parameters, therefore offering a flexible solution. This differs from ML-SGP4, where only the output of the SGP4 propagator is corrected, as explored in prior research [7]. In the next section, we will delve into some numerical experiments to showcase the use of ML-SGP4 and to highlight the comparison between ML-SGP4 and the ML-SGP4.